OpenStudy (anonymous):

A=1/2 bh b=17 & h=26

OpenStudy (amistre64):

and what are you spose to do with this?

OpenStudy (anonymous):

use the formula and values given to find the value of the reamaining variable?

OpenStudy (amistre64):

And where is it that you are having troubles with it?

OpenStudy (albert0898):

I agree with Amistre64. This shouldn't be a topic that you're having difficulty with. Just input the numbers. 1/2 * 17 * 26 Multiply it.

OpenStudy (anonymous):

i dont know where to start...i forgot how to do this

OpenStudy (albert0898):

Just multiply.

OpenStudy (anonymous):

i got 221

OpenStudy (albert0898):

Correct.

OpenStudy (amistre64):

you start by stating the formula A = 1/2 b h then you start replacing letters for their given values; since b = 17 A = 1/2 * 17 * h since h = 26, replace h by 26 A = 1/2 * 17 * 26 now since there are no more values to replace; finish it off by doing the math